OpenStudy (anonymous):

solve for x..... sin^2x+2sinx+1=0

OpenStudy (amistre64):

to factor the quadratic; its already a perfect square tto

OpenStudy (amistre64):

(sin(x)+1)^2 = 0 when sin(x) = -1

OpenStudy (anonymous):

don't you want to complete the square? lol

OpenStudy (anonymous):

idk how to do it with a sin though

OpenStudy (amistre64):

i would but..... aint nothing missing to it geometrically lol

OpenStudy (anonymous):

y = sin x y^2 - 2y + 1 = 0 y = 1 sin x = 1 x = pi/2 + n pi. where n is any integer, x is in radians.

OpenStudy (amistre64):

sin(x) is just another variable in this case; rename it if you want to i dunno; v^2

OpenStudy (anonymous):

okay you are all confusing me can someone just explain how to do it?

OpenStudy (amistre64):

Do you know how to solve it if it was: r^2 +2r +1 = 0 ?

OpenStudy (anonymous):

would it be sin = -1 ??

OpenStudy (anonymous):

or sinx= -1??

OpenStudy (amistre64):

yes; it would be sin(x) = -1

OpenStudy (anonymous):

yay!! thanks

OpenStudy (amistre64):

the extra bit is in knowing the values for x that make sin(x) = -1

OpenStudy (anonymous):

Sin x can be treated as a fancy variable or unknown. Just give it a simpler laber, say y. So if you replace sin x with y you have an prdinaty quadratic equation.

OpenStudy (anonymous):

okay so how do i find those values?

OpenStudy (amistre64):

sin(x) = -1 ; everytime x = 270 degrees ; its just knowing trigonometry which is alot of memorization

OpenStudy (anonymous):

so for my answer should i put sinx=-1 and 270 degrees?

OpenStudy (amistre64):

so unless we have an interval between 2 values to stay in; the proper answer is: x = 270 + n(360) ; where n is any integer

OpenStudy (anonymous):

y^2 +2y + 1 = (y +1)^2 (had the wronf sign before) y = -1 so sin x = -1 x = (2n - 1/2)*pi

OpenStudy (anonymous):

If x is in degrees, then x = (270 + 360n) degrees, n any integer

OpenStudy (anonymous):

Ninety degrees is one of the basic angles for the trig functions. 270 degrees puts you on the negative y-axis. If you are at (0,-1), then x = 0, y - 1, and r = 1(degeneerate triangle). sin(theta) = y/r = -1.